THE B-X TRANSITION OF B^{18}O

\underline{A. ANTIC-JOVANOVIC}, V. BOJOVIC, D. PESIC and M. KUZMANOVIC, Faculty of Physical Chemistry, University of Belgrade, P.O. Box 137, 11001 Belgrade, Yugoslavia.

The B − X bands of two isotopic species ^{11}B^{18}O and ^{10}B^{18}O have been observed for the first time. The bands were developed in emission from a low-pressure arc running in an atmosphere of the mixture of argon and oxygen 18 enriched up to 90 atom %. Twelve bandheads between 220 and 300 nm have been measured and vibrationaly assigned. The bands 1-0, 0-0 and 0-1 have been recorded under high dispersion and rotationally analyzed. Using determined molecular constants, Franck-Condon factors have been calculated and compared with visually estimated intensities of the bands. Full analyses of the bands will be presented.